OpenStudy (erickmanzo):

Solve 3x - 4 ≤ 2 or 2x + 11 ≥ -1. {x | x ≤ 2} {x | -6 ≤ x ≤ 2} {all reals}

OpenStudy (erickmanzo):

how do you solve it?????

OpenStudy (robtobey2):

\[3 x-4\leq 2\]3x-4+4<= 2+4 3 x <= 6 \[x\leq 2\]

OpenStudy (erickmanzo):

thank you

OpenStudy (precal):

but that was only part of the solution

OpenStudy (precal):

2x+11>-1 subtract 11 to both sides 2x>-12 divide 2 to both sides x>-6

OpenStudy (precal):

second one is your solution